# AVTENet: Audio-Visual Transformer-based Ensemble Network Exploiting Multiple Experts for Video Deepfake Detection

### v1

@@ III Proposed Method
The natural auditory context provides distinctive, independent, and diagnostic information about the visual world, which directly impacts the perceptual experience of visual objects [38]. To combat the emerging threat of video deepfakes, we leverage both acoustic and visual information to detect forgeries in videos. We use the transformer-based model because its self-attention mechanism can catch inconsistencies or manipulations to identify video forgeries. Furthermore, we employ ensemble learning to exploit the fusion of complementary information, modality-specific patterns, and multi-modal features to improve the robustness of the detection system. Accordingly, we propose an audio-visual transformer-based ensemble network (AVTENet) that fully exploits acoustic and visual information to detect video forgeries by leveraging modality-specific cues and complementary insights of joint audio and visual information. Fig. 3 shows an overview of our proposed AVTENet model, which consists of three key networks, namely a video-only network (VN), an audio-only network (AN), and an audio-visual network (AVN), as well as a decision-making module (DM). The three key networks are different-modality transformer-based networks integrated with pre-trained models through self-supervised learning (SSL) or supervised learning (SL). Then, DM integrates the outputs of these three classifiers according to different fusion strategies. Given a test video $x$, the decision is made according to
@@ III-A Video-only Network (VN)
The VN module extracts relevant spatiotemporal features from video frames through its self-attention mechanism to capture spatiotemporal patterns and long-range dependencies, and then outputs a vector representation that captures essential visual information.
Inspired by the remarkable progress achieved by the video vision transformer ViViT [52] in video classification tasks, we seamlessly integrate it into the visual backbone of AVTENet. As shown in the top part of Fig. 3, we use a factorized encoder model that consists of two separate transformer encoders. The first is a spatial encoder, which delves into the interaction among tokens extracted from the same temporal index. The second is a temporal encoder that models interactions among tokens from different temporal indices. We first divide the input video into small segments, a series of patches is extracted from each segment in the form of a tubelet and passed through the spatial encoder to produce an encoding vector representing the segment. The output encoding vectors serve as input to the temporal transformer along with an additional classification (cls) token to extract the final representation of the entire video.
For VN training, a dataset $D^{v}=\{v_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $v_{i}$ denotes the video stream of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $v_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the ViViT model extracts the single-vector representation $E_{v}$ (the $cls$ token of the temporal transformer) of each training video clip. A linear layer is used as the classifier. The ViViT model is pre-trained on the Kinetics dataset [52]. During the VN training process, not only is the linear classification layer trained, but also the pre-trained ViViT model is fine-tuned. The model is trained with the binary cross-entropy loss defined as
@@
where $x_{v}$ is the video stream (i.e., the image sequence) of $x$. Note that in the inference phase of our ensemble model, the DM module utilizes either the single-vector representation or the predicted score corresponding to the “fake” of the video-only network.
@@ III-B Audio-only Network (AN)
Transformers perform incredibly well in audio processing as their self-attention mechanisms allow for capturing long-range dependencies in audio. In this work, we use the Audio Spectrogram Transformer (AST) [53] as the audio backbone of AVTENet. As shown at the bottom of Fig. 3, the AN module takes as input a spectrogram that is further divided into a sequence of 16x16 patches with an overlap of 6 in both the time and frequency dimensions. A linear projection layer transforms each patch into a 1D embedding. Since the patch sequence lacks input order information, trainable positional embeddings are incorporated into each patch embedding to enable capturing spatiotemporal structure from a 2D audio spectrogram. Additionally, a classification (cls) token is added to the sequence. The output cls embedding of the transformer encoder encapsulates information about characteristics that can be used to determine challenging and deceptive alternations in the audio.
For training AN, a dataset $D^{a}=\{a_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ denotes the audio track of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $a_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AST model is used to extract the single vector representation $E_{a}$ (the $cls$ token) of the mel-spectrogram of each training audio clip. A linear layer is used as the classifier. The AST model is pre-trained on Audioset [53]. During the training process of AN, the linear classification layer is trained, and the pre-trained AST model is fine-tuned. As with the video-only network, the cross-entropy loss is used to train the audio-only network.
@@
where $x_{a}$ is the audio track of $x$. As with the video-only network, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to “fake” of the audio-only network is used by the DM module.
@@ III-C Audio-Visual Network (AVN)
We use the self-supervised learning framework AV-HuBERT [54] as the audio-visual backbone of AVTENet. As shown in the middle of Fig. 3, the AVN module simultaneously uses the audio spectrogram and the video frame stack of the lip region extracted from the video as inputs to the lightweight modality-specific encoders. The frames of the lip region are passed through a ResNet-based feature extractor to extract relevant visual features. At the same time, the audio spectrogram is passed through the feed-forward network (FFN) to extract the acoustic features. The visual and acoustic features are then fused and fed to a shared transformer encoder to extract jointly contextualized audio-visual representations that encapsulate the correlation between acoustic and visual modalities. These audio-visual representations are further passed through a Multiscale Temporal Convolutional Network (MS-TCN) and a temporal pooling layer to produce a single-vector representation.
For training AVN, a dataset $D^{av}=\{a_{i},v_{i}^{lip},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ and $v_{i}^{lip}$, respectively, denote the audio track and the lip image sequence of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $x_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AV-HuBERT model, MSTCN network, and temporal pooling layer are used to extract the vector representation $E_{av}$ of the mel-spectrogram of the audio track and the lip-image sequence of each training video clip. A linear layer is used as the classifier. The AV-HuBERT model is pre-trained on the LRS3 dataset [55]. During AVN’s training process, not only MSTCN with linear classification layer is trained, but also the pre-trained AV-HuBERT model is fine-tuned. As with the video-only and audio-only networks, the cross-entropy loss is used to train the audio-visual network.
@@
where $x_{a}$ and $x_{v^{lip}}$ are the audio track and the lip-image sequence of $x$, respectively. As with the video-only and audio-only networks, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to “fake” of the audio-visual network is used by the DM module.
In addition to the above AV-HuBERT-based AVN setup, we also study another setup. The audio and video embeddings are extracted via AST (same as AN) and ViViT (same as VN). The corresponding $cls$ tokens are concatenated as $E_{av}$. The prediction is also conducted by Eq. (5), but using ($x_{a}$, $x_{v}$) instead of ($x_{a}$, $x_{v^{lip}}$) as input. An overview of the AVTENet model based on this AVN setting is shown in Fig. 5. Compared with the AVTENet model in Fig. 3, the only difference is that the AV-HuBERT-based AVN is replaced by the AST&ViViT-based AVN.
@@ III-D Decision-Making Module (DM):
As shown in Fig. 3, the DM module of AVTENet is used to integrate the outputs of VN, AN, and AVN to produce final predictions that indicate potential forgery in the video. In this work, we investigate several fusion strategies.
@@ III-D1 Majority Voting
As shown in Fig. 4(a) and Algorithm 1, the DM module outputs “fake” if at least two component classifiers (VN, AN, and AVN) output “fake”. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{mv}$. Note that no additional models need to be trained for majority voting fusion.
@@
III-D2 Average Score Fusion
As shown in Fig. 4(b), the DM module outputs “fake” if the average of the output scores of the three component classifiers (VN, AN, and AVN) exceeds a preset threshold. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{asf}$. Note that no additional models need to be trained for average score fusion.
III-D3 Score Fusion
As shown in Fig. 4(c), in score fusion, the output scores of the three component classifiers (VN, AN, and AVN) are concatenated as input to a linear layer trained on the same training data as the three component classifiers. When training the linear layer, the parameters of the three component classifiers are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{sf}$.
III-D4 Feature Fusion
As shown in Fig. 4(d), in feature fusion, the output representations of the penultimate layers of the three component classifiers (VN, AN, and AVN) are concatenated as input to a linear layer trained on the same training data as the three component classifiers. When training the linear layer, the parameters of the three component classifiers are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{ff}$.
@@ IV-A Dataset
Our experiments are conducted on the FakeAVCeleb dataset [18], an audio-visual dataset released in 2022 specifically designed for the deepfake detection task. It is based on a collection of 500 YouTube videos featuring 500 ethnically diverse celebrities from various fields, including music, film, sports, and politics, taking into account age, geography, race, and gender diversity. Four latest deepfake manipulation and synthetic speech generation techniques, including Faceswap [48], FSGAN [49], Wav2Lip [50], and Real-Time-Voice-Cloning (RTVC) [51], and their combinations were used to generate 19,500 fake video samples from these 500 videos. As a result, the dataset contains a total of 20,000 video samples. The videos corresponding to 70 celebrities are used for testing, and the videos corresponding to the remaining 430 celebrities are used for training. These video samples are divided into four categories according to whether the audio and video modalities are manipulated, namely RealVideo-RealAudio ($R_{v}R_{a}$), RealVideo-FakeAudio ($R_{v}F_{a}$), FakeVideo-RealAudio ($F_{v}R_{a}$) and FakeVideo-FakeAudio ($F_{v}F_{a}$), where each video is carefully labeled to indicate its authenticity and facilitate training and testing forgery detection models.
@@
IV-A1 Testing Sets
To conduct a comprehensive evaluation of various uni-modal (AN and VN) and multi-modal (AVN and AVTENet) detection methods using the FakeAVCeleb dataset, we used eight different test sets, including Testset-I, Testset-II, faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, RTVC, and wav2lip. Except for Testset-I and Testset-II, each test set contains the same set of 70 genuine videos of 70 subjects unseen in training and a different set of 70 fake videos generated using specific manipulation techniques. Testset-I and Testset-II are the main evaluation test sets, which also contain the same 70 genuine videos of 70 subjects unseen in training and 70 fake videos. However, Testset-I contains the same number of fake samples from each manipulation technique, while Testset-II contains an equal number of fake samples from the $R_{v}F_{a}$, $F_{v}F_{a}$ and $F_{v}R_{a}$ categories. Table I summarizes the manipulation modalities involved in each test set. Multiple test sets can provide an exhaustive understanding of the strengths and limitations of AVTENet and individual classifiers on fake video samples generated using different deepfake techniques.
@@ IV-A2 Training Sets for Different Models
In the experiments, the training data are video samples corresponding to 430 subjects in FakeAVCeleb. However, training uni-modal and multi-modal classifiers requires specific data settings, so we customize the dataset according to the requirements of each network. Table II shows the training data setting for each network according to the categorical labels of FakeAVCeleb.
@@
Video Network Fake $F_{v}F_{a}$
$F_{v}R_{a}$
Real $R_{v}F_{a}$
$R_{v}R_{a}$
Audio Network Fake $R_{v}F_{a}$
$F_{v}F_{a}$
Real $F_{v}R_{a}$
$R_{v}R_{a}$
Audio-Visual Network Fake $F_{v}F_{a}$
$F_{v}R_{a}$
$R_{v}F_{a}$
@@ IV-A3 Data Analysis and Preprocessing
Each classifier involves different data processing steps to ensure effective training and evaluation. The video-only network takes as input a short clip of video; therefore, we split the video into multiple short clips and extract a sequence of consecutive frames from each clip to feed into the network. To prepare the data for the audio-only network, we extract the audio track with a sampling rate of 16kHz from each video. Furthermore, we extract mel-spectrograms from these audios, which are then used as input to the audio-only network. These mel-spectral features are also used as one of the inputs to the audio-visual network. For the AV-HuBERT-based AVN, we extract the lip image frames from each video and pair them with the mel-spectral features as input to the network.

### v2

@@ III Proposed Method
The natural auditory context provides distinctive, independent, and diagnostic information about the visual world, which directly impacts the perceptual experience of visual objects [6][46]. To combat the emerging threat of video deepfakes, inspired from human cognition of multisensory information, we leverage both acoustic and visual information to detect forgeries in videos. We use the transformer-based model because its self-attention mechanism can catch inconsistencies or manipulations to identify video forgeries. Furthermore, we employ ensemble learning to exploit the fusion of complementary information, modality-specific patterns, and multimodal features to improve the robustness of the detection system [7]. Accordingly, we propose an AVTENet that fully exploits acoustic and visual information to detect video forgeries by leveraging modality-specific cues and complementary insights of joint audio and visual information. Fig. 3 shows an overview of our proposed AVTENet model, which consists of three key networks, namely a video-only network (VN), an audio-only network (AN), and an audio-visual network (AVN), as well as a decision-making module (DM). The three key networks are different-modality transformer-based networks integrated with pre-trained models through self-supervised learning (SSL) or supervised learning (SL). Subsequently, DM integrates the outputs of these three classifiers according to different fusion strategies. Given a test video $x$, the decision is made according to
@@
The AVTENet architecture is designed to effectively capture and utilize complementary information from both modalities. Three networks are used in our ensemble approach to leverage the strength of each network. Individual unimodal networks (AN and VN) specialize in processing information from specific modalities. AN specializes in analyzing audio data to extract valuable acoustic features that help detect acoustic manipulation, while VN focuses on processing video data to extract visual features that can reveal visual forgeries. Combining them into AVN allows for a more comprehensive analysis of the input and can capture complex relationships and dependencies between the two modalities. These multimodal features provide more reliable detection results and enhance robustness.
@@ III-A Video-only Network (VN)
The VN module extracts relevant spatiotemporal features from video frames through its self-attention mechanism to capture spatiotemporal patterns and long-range dependencies. It then outputs a vector representation that captures essential visual information.
Building on the remarkable progress achieved by the video vision transformer ViViT [17] in video classification tasks, we seamlessly integrate it into the visual backbone of AVTENet. As shown in Fig. 3, we use a factorized encoder model that consists of two separate transformer encoders. The first is a spatial encoder, which delves into the interaction among tokens extracted from the same temporal index. The second is a temporal encoder that models interactions among tokens from different temporal indices. We first divide the input video into small segments, a series of patches is extracted from each segment as a tubelet and passed through the spatial encoder to produce an encoding vector representing the segment. The output encoding vectors serve as input to the temporal transformer along with an additional classification (cls) token to extract the final representation of the entire video.
For VN training, a dataset $D^{v}=\{v_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $v_{i}$ denotes the video stream of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $v_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the ViViT model extracts the single-vector representation $E_{v}$ (the $cls$ token of the temporal transformer) of each training video clip. A linear layer is used as the classifier. The ViViT model is pre-trained on the Kinetics dataset  [64]. During the VN training process, the linear classification layer is trained, and the pre-trained ViViT model is fine-tuned. The model is trained with the binary cross-entropy loss defined as
@@
where $x_{v}$ is the video stream (i.e., the image sequence) of $x$. In the inference phase of our ensemble model, the DM module utilizes either the single-vector representation or the predicted score corresponding to the “fake” of the video-only network.
@@ III-B Audio-only Network (AN)
Transformers perform incredibly well in audio processing as their self-attention mechanisms allow for capturing long-range dependencies in audio. In this study, we use the audio spectrogram transformer (AST) [18] as the audio backbone of AVTENet. As shown in Fig. 3, the AN module takes as input a spectrogram that is further divided into a sequence of 16x16 patches with an overlap of six in both the time and frequency dimensions. A linear projection layer transforms each patch into a 1D embedding. Since the patch sequence lacks input order information, trainable positional embeddings are incorporated into each patch embedding to enable capturing spatiotemporal structure from a 2D audio spectrogram. Additionally, a classification (cls) token is added to the sequence. The output cls embedding of the transformer encoder encapsulates information about characteristics that can be used to determine challenging and deceptive alterations in the audio.
For training AN, a dataset $D^{a}=\{a_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ denotes the audio track of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $a_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AST model is used to extract the single vector representation $E_{a}$ (the $cls$ token) of the mel-spectrogram of each training audio clip. A linear layer is used as the classifier. The AST model is pre-trained on the Audioset [18]. During the training of AN, the linear classification layer is trained, and the pre-trained AST model is fine-tuned. As with the video-only network, the cross-entropy loss is used to train the audio-only network.
@@
where $x_{a}$ is the audio track of $x$. As with the video-only network, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to the “fake” of the audio-only network is used by the DM module.
Figure 4: Architecture of MS-TCN block.
@@ III-C Audio-Visual Network (AVN)
We use the self-supervised learning framework AV-HuBERT [19] as the audio-visual backbone of AVTENet. As shown in Fig. 3, the AVN module simultaneously uses the log filterbank energies and the video frame stack of the lip region extracted from the video as inputs to the lightweight modality-specific encoders. The frames of the lip region are passed through a ResNet-based feature extractor to extract relevant visual features. Simultaneously, the log filterbank energy features are passed through the feed-forward network (FFN) to extract the acoustic features. The visual and acoustic features are then fused and fed to a shared transformer encoder to extract jointly contextualized audio-visual representations that encapsulate the correlation between acoustic and visual modalities. These audio-visual representations are further passed through a multiscale temporal convolutional network (MS-TCN) and a temporal pooling layer to produce a single-vector representation. Fig. 4 shows the structural diagram of MS-TCN. The details can be found in [65] and [66].
For training AVN, a dataset $D^{av}=\{a_{i},v_{i}^{lip},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ and $v_{i}^{lip}$, respectively, denote the audio track and the lip image sequence of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $x_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AV-HuBERT model, MS-TCN network, and temporal pooling layer are used to extract the vector representation $E_{av}$ of the log filterbank energy features of the audio track and the lip-image sequence of each training video clip. A linear layer is used as the classifier. The AV-HuBERT model is pre-trained on the LRS3 dataset [67]. During the training process of AVN, both the MS-TCN with a linear classification layer and the pre-trained AV-HuBERT model are trained, with the latter being fine-tuned.
@@
where $x_{a}$ and $x_{v^{lip}}$ are the audio track and the lip-image sequence of $x$, respectively. As with the video-only and audio-only networks, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to the “fake” of the audio-visual network is used by the DM module.
In addition to the above AV-HuBERT-based AVN setup, we also study another setup. The audio and video embeddings are extracted via AST (same as AN) and ViViT (same as VN). The corresponding $cls$ tokens are concatenated as $E_{av}$. The prediction is also conducted by Eq. (5), but using ($x_{a}$, $x_{v}$) rather than ($x_{a}$, $x_{v^{lip}}$) as input. An overview of the AVTENet model based on this AVN setting is shown in Fig. 5. Compared with the AVTENet model in Fig. 3, the only difference is that the AV-HuBERT-based AVN is replaced by the AST&ViViT-based AVN.
@@ III-D Decision-Making Module (DM):
As shown in Fig. 3, the DM module of AVTENet is used to integrate the outputs of VN, AN, and AVN (trained separately as described previously), to produce final predictions that indicate potential forgery in the video. In this study, we investigate several fusion strategies.
@@ III-D1 Majority Voting
As shown in Fig. 6(a) and Algorithm 1, the DM module outputs “fake” if at least two of VN, AN, and AVN output “fake”. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{mv}$. No additional models need to be trained for majority voting fusion.
III-D2 Average Score Fusion
As shown in Fig. 6(b), the DM module outputs “fake” if the average of the output scores of VN, AN, and AVN exceeds a preset threshold. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{asf}$. No additional models need to be trained for average score fusion.
III-D3 Score Fusion
As shown in Fig. 6(c), in score fusion, the output scores of VN, AN, and AVN are concatenated as input to a linear layer trained on the same training data as the three component classifiers with the cross-entropy loss. When training the linear layer, the parameters of VN, AN, and AVN are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{sf}$.
III-D4 Feature Fusion
As shown in Fig. 6(d), in feature fusion, the output representations of the penultimate layers of VN, AN, and AVN are concatenated as input to a linear layer trained on the same training data as the three component classifiers with the cross-entropy loss. When training the linear layer, the parameters of VN, AN, and AVN are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{ff}$.
@@ IV-A Dataset
Our experiments are conducted on the FakeAVCeleb dataset [^1] We do not use the DFDC dataset for two reasons. First, the subject’s face in some videos may not be facing the camera, and our model needs to extract lip images. Second, its labeling lacks separate labels for acoustic and visual modalities in videos. [25]. As described in Sec. II-C, it contains 500 videos featuring 500 celebrities and 19,500 fake videos manipulated from these 500 videos. The videos corresponding to 70 celebrities are used for testing, and the videos corresponding to the remaining 430 celebrities are used for training. These videos are divided into four categories according to whether the audio and video modalities are manipulated, namely RealVideo-RealAudio ($R_{v}R_{a}$), RealVideo-FakeAudio ($R_{v}F_{a}$), FakeVideo-RealAudio ($F_{v}R_{a}$) and FakeVideo-FakeAudio ($F_{v}F_{a}$), where each video is carefully labeled to indicate its authenticity and facilitate training and testing forgery detection models. Fig. 7 shows a few samples of each category from the FakeAVCeleb dataset.
IV-A1 Testing Sets
To conduct a comprehensive evaluation of various unimodal (AN and VN) and multimodal (AVN and AVTENet) detection methods using the FakeAVCeleb dataset, we used eight different test sets, including Testset-I, Testset-II, faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, RTVC, and wav2lip. Except for Testset-I and Testset-II, each test set contains the same set of 70 genuine videos of 70 subjects unseen in training and a different set of 70 fake videos generated using specific manipulation techniques. Testset-I and Testset-II are the main evaluation test sets, which also contain the above 70 genuine videos. However, for the 70 fake videos, Testset-I contains the same number of fake samples from each manipulation technique, while Testset-II contains an equal number of fake samples from the $R_{v}F_{a}$, $F_{v}F_{a}$ and $F_{v}R_{a}$ categories. Table I summarizes the manipulation modalities involved in each test set. Multiple test sets can provide an exhaustive understanding of the strengths and limitations of AVTENet and individual classifiers on fake video samples generated using different deepfake techniques.
@@ IV-A2 Training Sets for Different Models
In the experiments, the training data are video samples corresponding to 430 subjects in FakeAVCeleb. However, training unimodal and multimodal classifiers requires specific data settings, so we customize the dataset according to the requirements of each network. Table II shows the training data setting for each network according to the categorical labels of FakeAVCeleb. Table III shows the statistics of the real and fake samples used to train each network.
@@
VN Fake $F_{v}F_{a}$, $F_{v}R_{a}$
Real $R_{v}F_{a}$, $R_{v}R_{a}$
AN Fake $R_{v}F_{a}$, $F_{v}F_{a}$
Real $F_{v}R_{a}$, $R_{v}R_{a}$
AVN Fake $F_{v}F_{a}$, $F_{v}R_{a}$, $R_{v}F_{a}$
@@ IV-A3 Data Analysis and Preprocessing
Each classifier involves different data processing steps to ensure effective training and evaluation. The video-only network takes as input a short clip of video; therefore, we split the video into multiple short clips and extract a sequence of consecutive frames from each clip to feed into the network. To prepare the data for the audio-only network, we extract the audio track with a sampling rate of 16kHz from each video. Furthermore, we extract mel-spectrograms from these audios, which are then used as input to the audio-only network. We also extract log filterbank energy features that are used as one of the inputs to the audio-visual network. For the AV-HuBERT-based AVN, we extract the lip image frames from each video and pair them with the corresponding melspectral features as input to the network.